A small Malaysian electronics company located in Cyberjaya has been making customized OEM WiFi AC1900 routers usíng OpenWRT for a large multinational company, Huawei since 2015. Excellent business growth droved the company to determine the price and hourly demand for its most popular premium AC1900 router. After a detailed study the company arrived atthe following equation: D 1300-p where p is the price in RM and D is the hourly demand Using creative labour and resources management, the company is able to maintain fixed cost at RM1,500h and variable costs at RM100h.Dengan menggunakan maklumat yang diberikan tentukan perkara berikut: Using the given information, determine the following: Apakah bilangan optimum router harus dihasilkan setiap jam? What is the optimal number of routers should be produced each hour? a. (4 markah I marks) b. Berapa banyak router yang harus dihasilkan untuk mencapai titik pulang modal? How many routers should be produced to reach the breakeven point? (4 markah marks)
Answer #1

Answer

A) Given Demand function

D V1300-p

So D^2=1300-p

P=1300-D^2

So TR = D*P=D*(1300-D^2)=1300D-D^3

TC=1500+100=RM1600/Hr

For Profit S=TR-TC=1300D-D^3-1600

For Optimal condition dS/dD=0

1300-3D^2=0

D=21 units

B) For breakeven

TR=TC

1300D-D^3=1600

so solving for D we get

D= 2 units or 36 units

since we need to reach at breakeven so we will choose D=2 units